1. A random individual is chosen and given the test. Please show your work and find: a. P(person has condition | test is positive), b. P(person has condition | test is negative)

The conditions are as follows: Healthy people get a false positive 10% of the time

People with the condition get a false negative 10% of the time

5% of people actually have the condition.
Answer #1

a)P( person has condition | test is positive) =P(condition and tested positive)/P(tested positive)

=0.05*0.95/(0.05*0.95+0.95*0.1)=0.3333

b) P(person has condition | test is negative) =P(condition and tested negative)/P(tested negative)

=0.05*0.05/(0.05*0.05+0.95*0.9)=0.0029